OpenStudy (anonymous):

You are selling tickets for a high school basketball game. Student tickets cost $3 and general admission tickets cost $5. You sell 350 tickets and collect $1,450. How many student tickets did you sell?

OpenStudy (allank):

Fist, let the number of student tickets sold be, say, (s) So the number of general admission tickets will be (350-s) We can now get the money involved: 3(s)+5(350-s)=1450 Are you okay so far?

OpenStudy (anonymous):

im really dyslexic so i might be slow

OpenStudy (anonymous):

what i got was 200 but i dont think its right

OpenStudy (allank):

Well, I haven't calculated the final answer, but I can guide you to it. First, are you comfortable with my first statements up there?

OpenStudy (anonymous):

yea i do

OpenStudy (allank):

Good. 3(s)+5(350-s)=1450 So this becomes: 3s+17500-5s=1450 you okay with that?

OpenStudy (anonymous):

got it :) so far

OpenStudy (allank):

Excellent. So put like terms together: 3s-5s=1450-17500 Which becomes: -2s=-16050 right?

OpenStudy (anonymous):

right

OpenStudy (allank):

awesome. now divide both sides by -2 to get the value of s. what do you get?

OpenStudy (anonymous):

wait what?

OpenStudy (allank):

just a sec @primeavlqueen22 ...

OpenStudy (anonymous):

ok :)

OpenStudy (allank):

3(s)+5(350-s)=1450 3s+1750-5s=1450 -2s=1450-1750 -2s=-300 made a miscalculation

OpenStudy (allank):

so, like I said, we need to have s alone on the left side. to do that, we divide both the left and right sides by -2

OpenStudy (anonymous):

oooh ok i get thank you SO much your a lifesaver

OpenStudy (allank):

you're welcome :)